### v1

# Blowing up Dirac’s theorem

Abstract.
We show that every graph $G$ on $n$ vertices with $\delta(G)\geq(1/2+\varepsilon)n$ is spanned by a complete blow-up of a cycle with clusters of quasi uniform size $\Omega(\log n)$. The proof is based on a recently introduced approach for finding vertex-spanning substructures via blow-up covers.
1. Introduction
A fundamental problem in extremal combinatorics is to determine the least number of edges that guarantees a copy of a graph $F$ in an $n$-vertex host graph $G$. Erdős and Stone [5] proved that host graphs whose density slightly exceeds the Turán density $\tau(F)$, meaning $e(G)\geq(\tau(F)+\varepsilon)\binom{n}{2}$, even contain a blow-up of $F$. Formally, such a blow-up is defined by replacing the vertices of $F$ with vertex-disjoint clusters and each edge with a complete bipartite graph whose parts are the corresponding clusters. The arguments of Erdős and Stone show that we can find blow-ups of $F$ with clusters of size $\Omega(\log n)$, which is optimal as witnessed by dense random graphs.
It is natural, then, to ask if this ‘blow-up’ phenomenon extends to other types of structures than copies of a fixed graph $F$. We study this question for vertex-spanning subgraphs. In this setting, one typically replaces conditions on the number of edges with conditions on the minimum degree. For instance, the classic theorem of Dirac [3] states that every graph $G$ on $n\geq 3$ vertices with $\delta(G)\geq n/2$ contains a Hamilton cycle. In analogy to the Erdős–Stone Theorem, Böttcher, Schacht and Taraz [1] proved that a minimum degree of $\delta(G)\geq(1/2+\varepsilon)n$ guarantees a vertex-spanning blow-up of a cycle with clusters of order $\Omega(1)$. We extend this with the following blown-up version of Dirac’s theorem.
Theorem 1.1.
For every $\varepsilon,\eta>0$, there are $c>0$ and $n_{0}$ such that every graph $G$ on $n\geq n_{0}$ vertices with $\delta(G)\geq(1/2+\varepsilon)n$ contains a vertex-spanning blow-up of a cycle where each cluster has size in the range of $(1\pm\eta)c\log n$.
This result is related to recent work of the authors [7] for the analogous problem in hypergraphs and follows a similar proof strategy. However, unlike in the hypergraph setting, where the proof techniques limit us to clusters of $\operatorname{poly}(\log\log n)$ vertices, we are able to achieve asymptotically optimal cluster sizes in Theorem 1.1. It would be interesting to understand whether this can be extended to more complex structures such as powers of cycles.
2. Proof of the main result
In the following, we present two intermediate results, Lemmata 2.1 and 2.2, from which we then derive Theorem 1.1.
Let us begin by introducing some terminology. Consider a set family $\mathcal{V}$, whose elements are assumed to be pairwise disjoint by convention of this paper. A set $X$ is $\mathcal{V}$-partite if it has at most one vertex in each part of $\mathcal{V}$. We say that $\mathcal{V}$ is $m$-balanced if $|V|=m$ for every $V\in\mathcal{V}$. Similarly, $\mathcal{V}$ is $(1\pm\eta)m$-balanced if $(1-\eta)m\leq|V|\leq(1+\eta)m$ for every $V\in\mathcal{V}$. Moreover, $\mathcal{V}$ is quasi $(1\pm\eta)m$-balanced if there is a singleton $V^{\ast}\in\mathcal{V}$ and $\mathcal{V}\smallsetminus\{V^{\ast}\}$ is $(1\pm\eta)m$-balanced. Given a graph $R$ and a set family $\mathcal{V}=\{V_{v}\}$ indexed by $v\in V(R)$, the $\mathcal{V}$-blow-up is the blow-up obtained by replacing each $v$ with the set $V_{v}$. We denote it by $R(\mathcal{V})$, and say that $R$ is the reduced graph of the blow-up. Some of the constant hierarchies are expressed in standard $\gg$-notation. To be precise, we write $y\gg x$ to mean that for any $y\in(0,1]$ there exists an $x_{0}\in(0,1)$ such that for all $x_{0}\geq x$ the subsequent statements hold. Hierarchies with more constants are defined in a similar way and are to be read from left to right following the order that the constants are chosen. Moreover, we tend to ignore rounding errors if the context allows it.
Our first lemma states that we can cover any Dirac-type graph with vertex-disjoint blow-ups whose reduced graphs inherit the minimum degree up to a small error.
Lemma 2.1 (Simple blow-up cover).
Let $\varepsilon,1/s\gg c,\eta\gg 1/n$ and $m=c\log n$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Then the vertices of $G$ are covered by vertex-disjoint quasi $(1\pm\eta)m$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{t}(\mathcal{V}_{t})\subseteq G$ such that each reduced graph $R_{i}$ has $s$ vertices and satisfies $\delta(R_{i})\geq(1/2+\varepsilon/2)s$.
We note that Lemma 2.1 is a (quantitatively) refined version of a (qualitatively) more general result [7, Lemma 7.1]. The second lemma allows us to find a short chain of blow-ups, which connects two other clusters.
Lemma 2.2 (Connecting blow-ups).
Let $\varepsilon,c_{1}\gg c_{2}\gg 1/n$, $m=c_{1}\log n$ and $m^{\prime}=c_{2}\log n$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Let $\{U,V,W\}$ be a partition of $V(G)$ with $|U|=|V|=m$. Then there are $m^{\prime}$-sets $U^{\prime}\subseteq U$, $V^{\prime}\subseteq V$ and $W^{\prime}\subseteq W$ such that $G$ contains the complete bipartite graphs $K(U^{\prime},W^{\prime})$ and $K(V^{\prime},W^{\prime})$.
The proofs of Lemmata 2.1 and 2.2 can be found in section 4 and section 6, respectively. We continue by deriving our main result.
Proof of Theorem 1.1.
Introduce constants $s,c_{1},\eta,c_{2}$ with $\varepsilon\gg 1/s\gg c_{1},\eta\gg c_{2}\gg 1/n$ such that $\ell=c_{1}/c_{2}$ is an integer. (Note that the assumption $\varepsilon\gg\eta$ only gives a stronger outcome.) Set $m_{1}=c_{1}\log n$ and $m_{2}=c_{2}\log n$. Now consider an $n$-vertex graph $G$ with $\delta(G)\geq(1/2+\varepsilon)n$. Our goal is to find a $(1\pm 4\eta)m_{2}$-blow-up of a cycle, which allows us to conclude with $4\eta,c_{2}$ playing the rôles of $\eta,c$.
By Lemma 2.1, the vertices of $G$ are covered by vertex-disjoint quasi-$(1\pm\eta)m_{1}$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{t}(\mathcal{V}_{t})\subseteq G$ such that each reduced graph $R_{i}$ has order $s+1$ and satisfies $\delta(R_{i})\geq(1/2+\varepsilon/2)(s+1)$. We shall find a spanning blow-up of a cycle in each of the blow-ups $R_{i}(\mathcal{V}_{i})$. Fix $1\leq i\leq t$, and let $V^{\ast}\in\mathcal{V}_{i}$ be the (unique) singleton cluster of $\mathcal{V}_{i}$ with $V^{\ast}=V_{v^{\ast}}$ for some $v^{\ast}\in V(R_{i})$. The graph $R^{\prime}_{i}=R_{i}-v^{\ast}$ has $s$ vertices, and $\delta(R^{\prime}_{i})\geq(1/2+\varepsilon/2)(s+1)-1\geq s/2$. By Dirac’s theorem, $R^{\prime}_{i}$ contains a Hamilton cycle $C^{\prime}_{i}$, say with vertices $c_{1},\dotsc,c_{s-1}$. Using again that $\delta(R^{\prime}_{i})\geq s/2$ together with the pigeonhole principle, we find $1\leq j<s$ such that both $c_{j}$ and $c_{j+2}$ are neighbours of $v^{\ast}$ (index computations modulo $s$). Let $\mathcal{V}^{\prime}_{i}$ be obtained from $\mathcal{V}_{i}$ by adding the vertex of $V^{\ast}$ to $V_{c_{j}}$ (keeping the name for convenience) and then removing $V^{\ast}$ from the family. Note then that $R^{\prime}_{i}(\mathcal{V}^{\prime}_{i})\subseteq G$ is a $(1\pm 2\eta)m_{1}$-balanced blow-up, where $R^{\prime}_{i}$ has $s$ vertices and a Hamilton cycle $C_{i}$. Let $\mathcal{V}^{\prime}$ be the union of the families $\mathcal{V}^{\prime}_{i}$ defined in this way over all $1\leq i\leq t$, and note that $\mathcal{V}^{\prime}$ is $(1\pm 2\eta)m_{1}$-balanced with $ts$ clusters.
Now we can build our spanning blown-up up cycle greedily, by merging two cycles at a time using Lemma 2.2. More precisely, we will construct such connections, in an iterative fashion, joining two blow-ups at a time using an $m_{2}$-balanced blow-up $P_{i}(\mathcal{K}_{i})$ where $P_{i}$ is path of order $3$. Given $1\leq i\leq t$, let us describe the $i$th step of the construction, where we join the blow-ups $\mathcal{V}^{\prime}_{i}$ and $\mathcal{V}^{\prime}_{i+1}$ respectively, with corresponding reduced graphs $R^{\prime}_{i},R^{\prime}_{i+1}$ (indices taken modulo $t$). We assume that in previous steps corresponding to $1\leq j<i$, we have already built connecting $m_{2}$-sized blow-ups with families $\mathcal{K}_{1},\dotsc,\mathcal{K}_{j}$ of three clusters each. Let $K=\bigcup_{1\leq j<i}\bigcup\mathcal{K}_{j}$ be the union of the vertices of all those clusters. Since $t\leq n/m_{1}$, we have that
$|K|\leq 3tm_{2}\leq 3\frac{n}{c_{1}\log n}c_{2}\log n=\frac{3}{\ell}n\leq\frac{\eta\varepsilon}{8}n.$
Suppose that $C_{i}=c_{1}^{i}\dotsb c_{s-1}^{i}$ and $C_{i+1}=c_{1}^{i+1}\dotsb c_{s-1}^{i+1}$ are two Hamilton cycles in $R^{\prime}_{i},R^{\prime}_{i+1}$, respectively. Let $V^{i}_{s-1}\in\mathcal{V}^{\prime}_{i}$ and $V^{i+1}_{1}\in\mathcal{V}^{\prime}_{i+1}$ be the clusters corresponding to vertices $c^{i}_{s-1}\in V(R^{\prime}_{i})$ and $c^{i+1}_{1}\in V(R^{\prime}_{2})$, respectively.
We define a set of vertices which we want to avoid when building $\mathcal{K}_{i}$. Let $U\subseteq V(G)$ contain the vertices of clusters $V\in\mathcal{V}^{\prime}$ with $|V\cap K|\geq\eta m_{1}$. Note that there are at most $|K|/(\eta m_{1})$ such clusters $V$. So
$|U|\leq(1+2\eta)m_{1}\frac{|K|}{\eta m_{1}}\leq 2\frac{|K|}{\eta}\leq\frac{\varepsilon}{4}n,$
which gives $|U\cup K|\leq\varepsilon n/4+\eta\varepsilon n/8\leq\varepsilon n/2$. This implies that $G\smallsetminus(U\cup K)$ has minimum degree at least $(1/2+\varepsilon/2)n$. Apply Lemma 2.2 with $(G\smallsetminus(U\cup K))\cup V^{1}_{s-1}\cup V^{2}_{1}$, $V^{1}_{s-1}$, $V^{2}_{1}$, $m_{1}$, $m_{2}$ playing the rôles of $G$, $U$, $V$, $m$, $m^{\prime}$ respectively. This yields $m_{2}$-sets $W_{1}\subseteq V^{1}_{s-1}$, $W_{2}\subseteq V^{2}_{1}$, and $W_{3}$, so that the complete bipartite graphs $K(W_{1},W_{3})$ and $K(W_{2},W_{3})$ are contained in $G$. We finish by setting $\mathcal{K}_{i}=\{W_{1},W_{2},W_{3}\}$ and let $P_{i}$ be the corresponding $3$-vertex path. By construction, at the end of this process we used at most $\eta m_{1}$ connecting vertices in each of the clusters of $\mathcal{V}^{\prime}$.
Set $\mathcal{K}=\bigcup_{1\leq i\leq t}\mathcal{K}_{i}$, and let $\mathcal{V}^{\prime\prime}$ be the set family obtained from $\mathcal{V}^{\prime}$ after removing the vertices of $\mathcal{K}$ from each cluster. By construction, $\mathcal{K}$ is $m_{2}$-balanced and $\mathcal{V}^{\prime\prime}$ is $(1\pm 3\eta)m_{1}$-balanced. Moreover, there is a cycle $C^{\prime}$ on $t(s-1)+t=ts$ vertices such that we can index the family $\mathcal{K}\cup\mathcal{V}^{\prime\prime}$ by $V(C^{\prime})$, so that $C^{\prime}(\mathcal{V}^{\prime\prime}\cup\mathcal{K})$ is a spanning blow-up of a cycle in $G$, with all clusters of size at least $m_{2}$.
[…]
3. Tools
For the proofs of Lemmata 4.1 and 2.2, we require a few basic results from probability and extremal hypergraph theory. An $s$-uniform hypergraph (or $s$-graph for short) is a pair $P=(V,E)$ where every $e\in E$ is an $s$-element subset of $V$. The minimum degree $\delta_{1}(P)$ of an $s$-graph $P=(V,E)$ is the minimum of $\deg_{P}(v)=|\{e\in E\colon v\in e\}|$ over all $v\in V$.
Degree inheritance
A well-known probabilistic argument shows that relative minimum degree conditions are preserved with high probability when taking random induced subgraphs. Given a Dirac-type graph $G$, we use an $s$-uniform hypergraph $P$, called the property graph, which tracks the induced subgraphs that the minimum degree condition.
Lemma 3.1 (Degree inheritance).
Let $\varepsilon\gg 1/s\gg 1/n$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Let $P$ be the $s$-uniform hypergraph on the vertex set $V(G)$, where an $s$-set $S$ belongs to $P$ if and only if the induced graph $G[S]$ has minimum degree at least $(1/2+\varepsilon/2)s$. Then $\delta_{1}(P)\geq(1-e^{-\sqrt{s}})\tbinom{n-1}{s-1}$.
The proof of Lemma 3.1 follows from a concentration analysis. For sake of completeness, the details are presented in section A.
Blow-ups
Our approach relies on the existence of complete blow-ups with clusters of polylogarithmic size in dense hypergraphs, as guaranteed by a classic result of Erdős [4]. We use a refinement of this result for graphs by Nikiforov [8, 9], which allows us to find blow-ups with clusters of logarithmic size, not just polylogarithmic. The following is an immediate corollary of [8, Theorem 1].
Lemma 3.2 (Graph blow-ups).
Let $G$ be an $n$-vertex graph, and let $F$ be an $s$-vertex graph. Suppose that $G$ contains at least $\varepsilon n^{s}$ copies of $F$. Then $G$ contains a blow-up of $F$ with all clusters of size $\lfloor\varepsilon^{r^{2}}\log n\rfloor$.
4. Simple blow-up covers
The proof of Lemma 2.1 is adapted from our work on hypergraphs [7, §7]. We obtain better bounds on the cluster sizes by applying Lemma 3.2 instead of the related result for hypergraphs by Erdős [4]. Let us start with the following result, which has appeared (with constant cluster size) in the setting of perfect tilings [6, Corollary 10.2].
Lemma 4.1 (Almost blow-up cover).
Let $\varepsilon\gg 1/s\gg\eta\gg c\gg 1/n$ and $m=\lfloor c\log n\rfloor$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Then all but at most $\eta n$ vertices of $G$ may be covered with vertex-disjoint $m$-balanced blow-ups $\mathcal{V}_{1}(R_{1}),\dots,\mathcal{V}_{t}(R_{t})\subseteq G$ such that each reduced graph $R_{i}$ has $s$ vertices and satisfies $\delta(R_{i})\geq(1/2+\varepsilon/2)s$.
The next result allows us to pick up leftover vertices.
Lemma 4.2 (Rooted blow-ups).
Let $\varepsilon\gg 1/s\gg c\gg n$, and let $n^{1/2}\leq m_{1}\leq(\varepsilon/4)n$ and $m_{2}=c\log m_{1}$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Let $V\subseteq V(G)$ be an $m_{1}$-set. Then $G$ contains an $m_{2}$-balanced blow-up $T(\mathcal{W})$ such that $T$ has order $s$ and $\delta(T)\geq(1/2+\varepsilon/2)s$. Moreover, there is $W\in\mathcal{W}$ with $W\subseteq V$ and $W^{\prime}\cap V=\varnothing$ for all $W^{\prime}\in\mathcal{W}$ distinct from $W$.
We defer proofs of Lemmata 4.1 and 4.2 to section 5 and section 6 and continue to show the main result of this section.
Proof of Lemma 2.1 ( (Simple blow-up cover).).
Introduce $c_{1},c_{2},c_{3}$ with $c_{1},\eta\gg c_{2}\gg c_{3}\gg 1/n$. Let $m_{i}=c_{i}\log n$, for $i\in\{1,2,3\}$. By Lemma 4.1, all but at most $\eta n$ vertices of $G$ may be covered with pairwise vertex-disjoint $m_{1}$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{p}(\mathcal{V}_{p})\subseteq G$, where each $R_{i}$ has order $s$ and $\delta(G_{i})\geq(1/2+\varepsilon/2)s$. The uncovered vertices, denoted by $W$, are added onto blow-ups in two steps.
We first cover most vertices of $W$ as follows. While $|W|\geq\eta n/m_{1}$, iteratively apply Lemma 4.2 (with $|W|$ playing the rôle of $m_{1}$) to find vertex-disjoint blow-ups $T^{1}(\mathcal{W}_{1}),\dots,T^{p}(\mathcal{W}_{q})$ such that each $\mathcal{W}_{i}$ is an $m_{2}$-balanced set family with exactly one cluster contained in $W$. Each blow-up covers $(s-1)m_{2}$ vertices of $G-W$ and $m_{2}$ vertices of $W$. So the process requires at most $(s-1)|W|/m_{2}\leq n/m_{2}$ vertices of $G-W$, which means that we can greedily choose these blow-ups to be pairwise vertex-disjoint. Moreover, since $(s-1)m_{2}\leq\eta m_{1}$, we can also ensure that no cluster of $\mathcal{V}_{1}\cup\dots\cup\mathcal{V}_{p}$ is intersected in more than $(\eta/8)m_{1}$ vertices. Denote the remaining uncovered vertices by $U\subseteq W$.
To finish the covering, we apply Lemma 4.2 again to add each $u\in U$ onto a quasi $m_{3}$-balanced $\operatorname{\mathsf{P}}$-blow-up $S^{u}(\mathcal{U}_{u})$, where $S^{u}$ has order $s$ and $\{u\}\in\mathcal{U}_{u}$. (The requirements of the lemma can be met by replacing $u$ with $\varepsilon n$ copies of itself, which are dropped afterwards.) Note that this blow-up covers $(s-1)m_{3}$ vertices of $G-U$ and $1$ vertex of $U$. Since $(s-1)m_{3}|U|\leq\eta n$, we can select the blow-ups $S^{u}(\mathcal{U}_{u})$ ($u\in W$) to be pairwise vertex-disjoint. Moreover, since $(s-1)m_{3}\leq\eta m_{2}$, we may ensure that any cluster of $\mathcal{V}_{1},\dots,\mathcal{V}_{p},\mathcal{W}_{1}\cup\dots\cup\mathcal{W}_{q}$ is intersected in at most $(\eta/8)m_{2}$ vertices.
We delete the vertices of each $\mathcal{W}_{i}$ from each $\mathcal{V}_{j}$, keeping the names for convenience. Afterwards, we delete the vertices from $\mathcal{U}_{\ell}$ from each of $\mathcal{W}_{i}$ and $\mathcal{V}_{j}$, keeping again the names. To finish, we need to do two things: since the clusters of the families $\mathcal{W}_{i}$ and $\mathcal{V}_{i}$ are still much larger compared to $\mathcal{U}_{i}$, we need to split each of those clusters to make them smaller. Secondly, to obtain quasi balanced blow-ups, we need to create singleton clusters in those blow-ups. We address both of those issues as follows. First, we split each $\mathcal{W}_{i}$ and each $\mathcal{V}_{j}$ into $(1\pm\eta/2)((s-1)m_{3})$-balanced families. (This can be done greedily.) then, we partition each of these $(1\pm\eta/2)((s-1)m_{3})$-balanced families into $s$ quasi $(1\pm\eta)m_{3}$-balanced families. To do this, assuming the family is $\{X_{1},\dotsc,X_{s}\}$, can partition each $X_{i}$ into $s$ sets, one of size $1$, and the other ones of size as equal as possible, that is, of size $(1\pm\eta)m_{3}$. Then we form $s$ new families by including in each one exactly one $1$-sized set, and $s-1$ many $(1\pm\eta)m_{3}$-sized sets. Hence every family is now quasi $(1\pm\eta)m_{3}$-balanced, and we can finish with $c=c_{3}$. ∎
5. Almost blow-up covers
Here we show Lemma 4.1. We require the following notion of (weak) quasirandomness for $s$-uniform hypergraphs. For an $s$-graph $P$, the density of a tuple $(V_{1},\dots,V_{k})$ of pairwise disjoint vertex sets is
$d_{P}(V_{1},\dots,V_{k})=\frac{e_{P}(V_{1},\dots,V_{k})}{|V_{1}|\dotsb|V_{k}|}.$
We say that $(V_{1},\dots,V_{k})$ is $(\varrho,d)$-lower-regular if for $d=e_{P}(V_{1},\dots,V_{k})$ and all choices of $X_{1}\subseteq V_{1},$ $\dots$,… , $X_{k}\subseteq V_{k}$ satisfying $|X_{1}|\geq\varrho|V_{1}|,\,\dots,\,|X_{k}|\geq\varrho|V_{k}|$, we have $d_{P}(X_{1},\dots,X_{k})\geq d-\varrho$. Denote by $\mathcal{Q}(s,m,\varrho,d)$ the set of $(\varrho,d^{\prime})$-lower-regular $s$-tuples which are $m^{\prime}$-balanced, with $m^{\prime}\geq m$, and $d^{\prime}\geq d$.
Our next lemma shows that one can partition most of the vertices of a hypergraph with large enough minimum degree into balanced lower-regular tuples. This is a straightforward consequence of the Weak Hypergraph Regularity Lemma [6, Theorem B.2]. A simpler, alternative proof of this fact appears in our related work on hypergraphs [7, §7.3]. For sake of completeness, the details are presented in section B.
For an $s$-graph $P$ and a family $\mathcal{F}$ of $s$-graphs, an $\mathcal{F}$-tiling in $P$ is a set of pairwise vertex-disjoint $k$-graphs $F_{1},\dots,F_{\ell}\subseteq P$ with $F_{1},\dots,F_{\ell}\in\mathcal{F}$.
Lemma 5.1 (Almost perfect tiling).
Let $1/s\gg\eta,\varrho\gg\alpha\gg 1/n$ and $m=\alpha n$. Then every $n$-vertex $s$-graph $P$ with $\delta_{1}(P)\geq\left(1-1/s^{2}\right)\binom{n-1}{s-1}$ contains a $\mathcal{Q}(s,m,\varrho,2\eta)$-tiling that covers all but $\eta n$ vertices.
Given this, we can easily derive Lemma 4.1, the main result of this section.
Proof of Lemma 4.1.
Introduce $\alpha$ with $\eta\gg\alpha\gg c$, and set $\varrho=\eta/2$. We let $P$ be the auxiliary $s$-uniform hypergraph on the vertex set $V(G)$, where an $s$-set $S$ belongs to $P$ if and only if the induced graph $G[S]$ has minimum degree at least $(1/2+\varepsilon/2)s$. By Lemma 3.1, we have that $\delta_{1}(P)\geq(1-e^{-\sqrt{s}})\binom{n-1}{s-1}\geq(1-1/s^{2})\binom{n-1}{s-1}$.
By Lemma 5.1, the $s$-graph $P$ contains a $\mathcal{Q}(s,\alpha n,\varrho,\eta)$-tiling that covers all but $\eta n$ vertices. Consider one of the $(\varrho,\eta)$-lower-regular $s$-tuples with $\alpha n$-balanced parts $\mathcal{V}=\{V_{1},\dots,V_{s}\}$. Note that each $\mathcal{V}$-partite edge of $P$ corresponds to some labelled $s$-vertex graph with minimum degree at least $(1/2+\varepsilon/2)s$. Note that there are at most $2^{s^{2}}$ different labelled $s$-vertex graphs. By the pigeonhole principle there is an $s$-vertex graph $R_{1}$ with $\delta(R_{1})\geq(1/2+\varepsilon/2)s$ such that at least $(\eta-\varrho)2^{-s^{2}}(\alpha n)^{s}$ edges of $P$ correspond to a labelled copy of $R_{1}$. By Lemma 3.2, we may find a subgraph $R_{1}(m)\subseteq G$ with $\mathcal{V}$-partite edges, which constitutes one of the desired blow-ups. We repeat this procedure as long as there are at least $(\eta-\varrho)(\varrho\alpha n)^{s}$ $\mathcal{V}$-partite unused edges, thus obtaining pairwise vertex-disjoint blow-ups $R_{1}(m),\dots,R_{\ell}(m)$. Since the tuple $(V_{1},\dots,V_{s})$ is $(\varrho,\eta)$-lower-regular, this procedure stops when all but at most $\varrho\alpha n$ vertices of $\mathcal{V}$ are used in each cluster of $\mathcal{V}$. After iterating this over all regular tuples, we have thus covered all but $\varrho n+\eta n\leq 2\eta n$ vertices, which allows us to conclude with $2\eta$ in place of $\eta$. ∎
6. Connecting and rooting blow-ups
In this section, we show Lemmata 2.2 and 4.2, starting with the former.
Proof of Lemma 2.2.
Note that every vertex in $U\cup V$ has at least $(1/2+\varepsilon/2)n$ neighbours in $W$. Let $W_{1}\subseteq W$ be the set of vertices of $W$ with at least $\varepsilon m/8$ neighbours in $U$. Counting edges between $U$ and $W$, we get $(1/2+\varepsilon/2)mn\leq|W_{U}|m+(n-|W_{U}|)\varepsilon m/8$, which gives
$|W_{U}|\geq\frac{1+\varepsilon-\varepsilon/4}{2-\varepsilon/4}n\geq\left(\frac{1}{2}+\frac{\varepsilon}{2}\right)n.$
Now let $W_{V}\subseteq W$ be the set of vertices of $W$ with at least $\varepsilon m/8$ neighbours in $V$. An analogous argument gives that $|W_{U}|\geq\left(1/2+\varepsilon/2\right)n$. Hence, the set $W^{\ast}=W_{U}\cap W_{V}$ has size at least $\varepsilon n$.
Set $t=m^{\prime}$. We claim that there are $t$-sets $W_{1}\subseteq U$, $W_{2}\subseteq V$ and $W_{3}\subseteq W$ such that $G$ contains the complete bipartite graphs $K(W_{1},W_{3})$ and $K(W_{2},W_{3})$. Suppose otherwise. So no two $t$-sets in $U$ and $V$ have more than $t-1$ common neighbours in $W^{\ast}$. Therefore
$\sum_{w\in W^{\ast}}\binom{\deg(w,U)}{t}\binom{\deg(w,V)}{t}\leq(t-1)\binom{m}{t}^{2}.$
On the other hand, we have
$\sum_{w\in W^{\ast}}\binom{\deg(w,U)}{t}\binom{\deg(w,V)}{t}\geq\varepsilon n\binom{\varepsilon m/8}{t}^{2}\geq\varepsilon n(\varepsilon/8)^{2t}\binom{m}{t}^{2},$
where we used $t\leq\varepsilon m/2$ in the last inequality. Combining both bounds, we get $\varepsilon n(\varepsilon/8)^{2t}\leq t$. Hence,
$\log(\varepsilon)+\log n-\log(8/\varepsilon)2t\leq\log t,$
which contradicts the choice of $t$. ∎
It remains to show how to obtain rooted blow-ups.
Proof of Lemma 4.2.
Let $G^{\prime}$ be the graph obtained from $G$ by removing the edges inside $V$. Since $|V|=m_{1}\leq\varepsilon n/4$ this barely affects the minimum degree. So we still have $\delta(G^{\prime})\geq(1/2+3\varepsilon/4)n$. Let $P$ be the $s$-uniform hypergraph where an $s$-set forms an edge if it has exactly one vertex in $V$ and the induced subgraph $G^{\prime}[S]$ has minimum degree at least $(1/2+\varepsilon/2)s$. By Lemma 3.1 and crude bounding, we have $e(P)\geq|V|n^{s-1}/2$. Consider a random partition of $V(G)\smallsetminus V$ by assigning each vertex uniformly to one of $s-1$ parts. Together with $V$, this gives a partition of $V(G)$ into $s$ parts. Let $P^{\prime}\subseteq P$ be the $s$-partite $s$-graph, which contains the crossing edges. By a first moment argument, it follows that with positive probability $e(P^{\prime})\geq|V|n^{s-1}k!/(2k^{k})$. Fix such a partition for the rest of the proof. Recall that there are at most $2^{s^{2}}$ labelled graphs on $s$ vertices. Let $R$ be the labelled graph that appears on the edges of $P^{\prime}$ most frequently, where the labels correspond to the parts. We pass to a subgraph $P^{\prime\prime}\subseteq P^{\prime}$ keeping only the $s$-sets $S$ where $G[S]$ is isomorphic to $R$. So $P^{\prime\prime}$ has at least $|V|n^{s-1}k!/(k^{k}2^{s^{2}+2})$ edges. Note that by our choice we can assume that for each $S\in P^{\prime\prime}$, $G[S]$ is isomorphic to $R$, $|S\cap V|=1$ and also $G[S\smallsetminus V]$ is isomorphic to the same $(s-1)$-graph $R^{\prime}$, no matter the choice of $S$.
Now let $1/s\gg\gamma_{1}\gg\gamma_{2}\gg c$, so that $e(P^{\prime\prime})\geq\gamma_{1}|V|n^{s-1}$. Let $\mathcal{R}^{\prime}$ be the set of ordered $(s-1)$-tuples in $G-V$, inducing a copy of $R^{\prime}$, which can be extended to at least $\gamma_{1}|V|/2$ copies of $R$ by adding a vertex of $V$. Since $e(P^{\prime\prime})\geq\gamma_{1}|V|n^{s-1}$, we have that $|\mathcal{R}^{\prime}|\geq\gamma_{1}n^{s-1}/2$. By Lemma 3.2, there is a blow-up $K\subseteq G-V$ of $R^{\prime}$ with all clusters of size $t=\gamma_{2}\log n$. Let $R^{\prime}_{1},\dotsc,R^{\prime}_{t}$ be vertex-disjoint copies of $R^{\prime}$ in the blow-up $K$. Form an auxiliary bipartite graph $B$ with vertex classes $V$ and $W=\{R^{\prime}_{1},\dotsc,R^{\prime}_{t}\}$ where $u\sim R^{\prime}_{i}$ if $R^{\prime}_{i}$ can be extended to a copy of $R$ by adding $u$. Then $B$ has at least $\gamma_{1}t|V|/2$ edges.
Note that to conclude it is enough to find a complete bipartite graph in $B$ with parts of size $c\ln n$. The existence of such graph follows from the usual bounds for the Zarankiewicz problem. To be precise, let $p=c\log n$ and suppose $B$ does not have a complete bipartite graph with two classes of size $p$. Then each $p$-set $X\subseteq W$ has at most $p-1$ common neighbours $N_{B}(X)$ in $V$. Thus, double-counting and convexity reveals
$(p-1)\binom{t}{p}$ $\geq\sum_{X\in\binom{W}{p}}|N_{B}(X)|=\sum_{v\in V}\binom{\deg_{B}(v)}{p}$
$\geq|V|\binom{e(B)/|V|}{p}\geq|V|\binom{\gamma_{1}t/2}{p}\geq(\gamma_{1}/4)^{p}|V|\binom{t}{p},$
where we used that $p=c\log n\leq\gamma_{2}\log n=\gamma_{1}t/4$ in the last inequality. Hence
$p\geq(\gamma_{1}/4)^{p}|V|\geq|V|n^{-1/4}\geq n^{1/4},$
where the second inequality follows from the (small enough) choice of $c$, and the last inequality follows from $|V|=m_{1}\geq n^{1/2}$. This contradicts the choice of $p$, and proves the result. ∎
Acknowledgements
Richard Lang was supported by the Ramón y Cajal programme (RYC2022-038372-I) and by grant PID2023-147202NB-I00 funded by MICIU/AEI/10.13039/501100011033. Nicolás Sanhueza-Matamala was supported by ANID-FONDECYT Iniciación Nº11220269 grant.
A. Inheritance
For a $k$-graph $G$ on $n$ vertices, we denote the relative degree of vertex $v\in V(G)$ by $\overline{\deg}_{G}(v)=\deg_{G}(v)/(n-1)$. The following instance of the inheritance principle immediately implies Lemma 3.1.
Lemma A.1 (Degree inheritance).
For every $\varepsilon>0$, there is $C>0$ such that the following holds for all $1\leq s\leq n$. Suppose $G$ is an $n$-vertex graph. Let $P$ be the $s$-graph on $V(G)$ with an edge $S$ whenever every vertex $v\in S$ satisfies $\overline{\deg}_{G[S]}(v)\geq\overline{\deg}_{G}(v)-\varepsilon$. Then $\delta_{1}(P)\geq\left(1-\exp({-Cs})\right)\tbinom{n-1}{s-1}$.
We show Lemma A.1 using the following standard tail bound. Recall that a hypergeometrically distributed random variable $X\sim\text{H}(N,K,n)$ describes the probability of $k$ successes in $n$ draws, without replacement, from a finite population of size $N$ that contains exactly $K$ objects with that feature, wherein each draw is either a success or a failure. Moreover, the expected value of $X$ is $\mathbb{E}(X)=nN/K$.
Lemma A.2 (Hoeffding’s inequality).
For $X\sim\text{H}(N,K,n)$ and any $0<\ell<nK/N$, we have ${\rm Pr}(|X-\mathbb{E}(X)|\leq\ell)\leq 2\exp\left(-2\ell^{2}/n\right)$.
Proof of Lemma A.1.
Set $V=V(G)$, and fix an arbitrary vertex $x\in V$. Let $S\subseteq V$ be drawn uniformly among all $s$-sets. We say that vertex $v\in V$ is bad for $S$ if $\overline{\deg}_{G[S\cup\{v\}]}(v)<\overline{\deg}_{G}(v)-\varepsilon$. Moreover, $S$ is good if it contains no vertex which is bad for $S$. So $P$ is the $s$-graph of good edges. It suffices to show that, conditioning on the event that $x\in S$, every vertex $v\in V$ is bad for $S$ with probability at most $\exp(-\Omega(s))$. Indeed, in this case,
$\Pr(\text{$S$ is good}\mid x\in S)$ $\leq\sum_{v\in V}\Pr(\text{$v$ is bad for $S$}\mid\{x,v\}\subseteq S)$
$\leq\sum_{v\in V}\Pr(\text{$v$ is bad for $S$}\mid x\in S)\Pr(v\in S)$
$=\exp({-\Omega(s)})\sum_{v\in V}\Pr(v\in S)=s\exp({-\Omega(s)})=\exp({-\Omega(s)}).$
To obtain the above bound, consider a vertex $v\in V$ with $\delta=\overline{\deg}_{G}(v)-\varepsilon$. We can assume that $\delta\geq 0$ as otherwise there is nothing to show. Let $q$ be $1$ if $xv\in E(G)$ and 00 otherwise. Let $X=\deg_{G[S\cup\{v\}]}(v)$ conditioning on $x\in S$, and note that $X\sim q+\text{H}(n-1,\deg_{G}(v)-q,s-1)$. So $\mathbb{E}(X)\geq(s-1)(\deg_{G}(v)-q)/(n-1)\geq(\delta+(3/4)\varepsilon)(s-1)$. It follows by Lemma A.2 applied with $s-1$, $(\varepsilon/4)(s-1)$ playing the rôles of $n$, $\ell$ that
${\rm Pr}\left(X<\delta(s-1)\right)$ $\leq{\rm Pr}\left(\mathbb{E}(X)-X\geq(\varepsilon/4)(s-1)\right)$
$\leq 2\exp\left(-\frac{2(\varepsilon/4)^{2}}{s}(s-1)^{2}\right)\leq\exp(-\Omega(s)).\qed$
B. Almost perfect tilings
In this section, we show Lemma 5.1 following the exposition of our work on hypergraphs [7, §7].
Regular tuples
At the heart of the proof is the fact that every (reasonably) large dense hypergraph admits a regular tuple of linear order.
Lemma B.1 (Regular tuple).
Let $1/s,\,d,\,\varrho,\nu\gg 1/m$ and $\alpha=\exp(-\varrho^{-2s})$. Let $P$ be an $m$-balanced $s$-partite $s$-graph with $e(P)\geq dm^{s}$. Then $P$ contains an $(\varrho,d)$-lower-regular $m_{1}$-balanced $s$-tuple, where $\alpha m\leq m_{1}$.
Proof.
[…]
Hypergraph matchings
We also require the following (crude) Dirac-type bound for hypergraph matchings due to Daykin and Häggkvist [2], whose short proof is included for sake of completeness.
Lemma B.2.
Let $1/s,\varepsilon\gg 1/n$ with $n$ divisible by $s$. Let $P$ be an $n$-vertex $s$-graph with $\delta_{1}(P)\geq(1-1/s+\varepsilon)\tbinom{n-1}{s-1}$. Then $P$ has a perfect matching.
Proof.
Set $m=n/s$. To begin, we randomly partition the vertices of $P$ into $k$ parts of size $m$. Let $P^{\prime}$ be the $n$-vertex $s$-partite $s$-graph obtained by keeping only the partite edges. Using standard concentration bounds (such as McDiarmid’s inequality), we see that $\delta_{1}(P^{\prime})\geq(1-1/s+\varepsilon/2)m^{s-1}$ with high probability. Fix such a partition. We shall prove that $P^{\prime}$ contains a perfect matching.
Consider a largest matching $M$ in $P^{\prime}$. We can assume that $M$ covers all but $s$ vertices of $P^{\prime}$. Otherwise just add some dummy edges to $M$. Let $v_{1},\dots,v_{s}$ be the vertices not covered by $M$, where $v_{i}$ is in cluster $V_{i}$ of $P^{\prime}$. Let $\sigma_{1},\dots,\sigma_{s}\in\mathcal{S}_{s}$ be arbitrary permutations with $\sigma_{i}(i)=i$. Let $X=(e_{1},\dots,e_{s-1})\in M^{s-1}$. For $\sigma_{i}$ and $1\leq j\leq s$ with $j\neq i$, let by $v_{\sigma_{i}}(j)$ be the vertex of $e_{\sigma(i)}$ in cluster $V_{j}$. Set $v_{\sigma_{i}}(i)=v_{i}$. We say that $X$ is good for index $i$, if $\{v_{\sigma_{i}}(1),\dots,v_{\sigma_{i}}(s)\}\in E(P^{\prime}).$ There are at most $m^{s-1}$ choices for $X$. Note that at most $\binom{s-1}{2}n^{s-2}<\varepsilon/2$ of theses choices contain a repeated matching edge. Moreover, for fixed $i$, there are at least $(1-1/s+\varepsilon/2)m^{s-1}$ good $X\in M^{s-1}$. (One for each $M$-partite neighbour of $v_{i}$.) So there is some $X$ without repeated matching edge, which is good for all $1\leq i\leq s$. Now we just have to choose $\pi_{1},\dots,\pi_{s}$ such that we may find a perfect matching between the vertices of $X$ and and $v_{1},\dots,v_{s}$. But this is an easy exercise. So we have a new matching $M^{\prime}$. If $M$ did not have dummy edges, we are done. Otherwise we have found a larger matching than $M$, which is a contradiction. ∎
Incrementalism
Now we are ready to show the main result of this section. The next lemma will turn a $\mathcal{Q}$-tiling into another $\mathcal{Q}$-tiling with smaller parts which covers more vertices, so it suffices to iterate until most vertices are covered. More precisely, we apply the lemma with $1/s\gg\mu\gg\alpha$ and $\eta=\mu/16$, so that $1-1/s^{2}\geq 1-1/s+\mu$. In each step, we turn a $\mathcal{Q}$-tiling into another $\mathcal{Q}$-tiling with smaller parts such that additional $(\nu\mu/16)n=\eta^{2}n$ vertices are covered. So we arrive at Lemma 5.1 after at most $\eta^{-2}$ steps.
Lemma B.3 (Regular tuple tiling increment).
Let $1/s,\,\mu,\,\eta,\,d,\,\varepsilon\gg\alpha\gg 1/n$ with $d\leq\mu/16$ and $\varepsilon\leq d/2$. Set $\gamma=\exp(-\varepsilon^{-2s})$ and $m=\alpha n$. Let $\mathcal{Q}_{1}=\mathcal{Q}(s,m,\varepsilon,{d})$ and $\mathcal{Q}_{2}=\mathcal{Q}(s,m^{\prime},\varepsilon^{\prime},d^{\prime})$, where $m^{\prime}=(\gamma\eta\mu^{2}/128)m$, $\varepsilon^{\prime}=2\varepsilon$, and $d^{\prime}=d/2$. Let $P$ be an $s$-graph on $n$ vertices with $\delta_{1}(P)\geq\left(1-1/s+\mu\right)\binom{n-1}{s-1}$. Suppose that $P$ contains a $\mathcal{Q}_{1}$-tiling $Q_{1}$ on $\lambda n$ vertices with $0\leq\lambda\leq 1-\eta$. Then $P$ contains a $\mathcal{Q}_{2}$-tiling on at least $(\lambda+\mu\eta/16)n$ vertices.
Before we come to the proof of Lemma B.3, let us derive Lemma 5.1.
Proof of Lemma 5.1.
Set $\mu=16\eta$, and note that $1-1/s^{2}\geq 1-1/s+\mu$. Set $t=(\mu\eta/16)^{-1}=\eta^{-2}$. For $0\leq i\leq t$, we set $d_{i}=(\mu/16)/2^{i}$, $\varepsilon_{i}=d_{i}/2$, $\gamma_{i}=\exp(-\varepsilon_{i}^{-2s})$, $\alpha_{0}=1$, $\alpha_{i+1}=\gamma_{i+1}\eta\mu^{2}/128$, $m_{i}=\alpha_{i}n$, $\lambda_{0}=0$ and $\lambda_{i+1}=\lambda_{i}+\eta^{2}$. We begin with $\mathcal{Q}_{1}=\mathcal{Q}(s,\alpha_{0}n,\varepsilon_{0},d_{0})$ and an empty $\mathcal{Q}_{1}$-tiling $Q_{1}$. We then apply Lemma B.3 iteratively until all but $\eta n$ vertices are covered. So at the beginning of step $i$, we have a $\mathcal{Q}_{1}$-tiling $Q_{1}$ on $\lambda_{i}n$ vertices in $P$, where $\mathcal{Q}_{1}=\mathcal{Q}(s,m_{i},\varepsilon_{i},d_{i})$. And at the end of step $i$, we have a $\mathcal{Q}_{2,i}$-tiling on at least $(\lambda_{i}+\mu\eta/16)n=(\lambda_{i}+\eta^{2})n=\lambda_{i+1}n$ vertices, where $\mathcal{Q}_{2,i}=\mathcal{Q}(s,m_{i+1},\varepsilon_{i+1},d_{i+1})$. Clearly, this process stops after at most $t$ steps. So we can finish with a $\mathcal{Q}(s,\alpha_{t}n,\varepsilon_{t},d_{t})$-tiling that covers all but $\eta n$ vertices. ∎
Let us briefly explain how we intend to prove Lemma B.3. Our plan is to extract additional regular tuples from the $(1-\lambda)n\geq\eta n$ vertices not covered by $Q_{1}$. We shall do this by using the assumption that the minimum degree is $\mu\binom{n-1}{s-1}$ above the threshold for a perfect matching. To track these gains, we define $\nu=\mu/16$ and $\beta=\nu\eta/2$. In a first step, we find a tiling $\mathcal{Q}_{\text{fresh}}$ that covers $4\beta\eta n$ additional vertices outside of $Q_{1}$ and some vertices inside of $Q_{1}$. In a second step, we ‘recycle’ the remainder of $Q_{1}$ using with another tiling $\mathcal{Q}_{\text{rec}}$ such that $\mathcal{Q}_{\text{fresh}}\cup\mathcal{Q}_{\text{rec}}$ together cover all but $2\beta\eta n$ vertices in $Q_{1}$. We note that the leftover comes from the fact that removing $\mathcal{Q}_{\text{fresh}}$ leaves parts of $Q_{1}$ a bit unbalanced. In any case, the relative difference $162\beta n=\mu\eta/$ 16 presents the share of additional covered vertices at the end of the argument.
Recall that the input part sizes are $m$ and $m^{\prime}$. Since the parts of $Q_{1}$ may be of different size, it will be convenient to use an intermediate part size. We therefore define $m_{1}=m$, $m_{2}=(\mu/8)m_{1}$ and $m_{3}=m^{\prime}=\gamma 2\beta m_{2}$. Our approach is then to refine the parts of $Q_{1}$, which have size at least $m_{1}$, into parts of uniform size $m_{2}$. We remark that $m_{2}$ is chosen small enough to take advantage of the excess minimum degree $\mu\binom{n-1}{s-1}$. Moreover, $m_{3}$ is chosen small enough to keep the number of lost vertices due to imbalancedness (discussed above) under $2\beta n$. Now come the details.
Proof of Lemma B.3.
We set $\nu=\mu/16$ and $\beta=\nu\eta/2$. Moreover, let $m_{1}=m$, $m_{2}=(\mu/8)m_{1}$ and $m_{3}=m^{\prime}=\gamma 2\beta m_{2}$. We begin by defining family $\mathcal{U}$ of disjoint sets in $V(P)$, as follows. For each $s$-tuple in $Q_{1}$ with parts $\{V_{1},\dotsc,V_{s}\}$, pick a maximal number of vertex-disjoint sets of size $m_{2}$ inside each $V_{i}$, and add all of them to $\mathcal{U}$. It follows the sets of $\mathcal{U}$ cover all but $m_{2}=(\mu/8)m_{1}\leq(\mu/8)|V_{i}|$ vertices in each $V_{i}$. Next, we extend $\mathcal{U}$ by subdividing the vertices outside of $Q_{1}$ into a maximal number of vertex-disjoint $m_{2}$-sets such that $r=|\mathcal{U}|$ is divisible by $s$. This implies that outside of $V(Q_{1})$, the family $\mathcal{U}$ covers all but at most $sm_{2}$ vertices. We remark that $1/s,\mu$ are much larger than $1/r$ by the choice of $\alpha$. Now, let $R$ be an $s$-graph with vertex set $\mathcal{U}$ and an edge $X$ if $P$ contains at least $4\nu m_{2}^{s}$ $X$-partite edges. So $R$ plays the rôle of a reduced graph in the context of a Regularity Lemma. It is therefore not surprising that $R$ inherits the minimum degree condition of $P$.
Claim B.4.
We have $\delta_{1}(R)\geq\left(1-1/s+\mu/2\right)\binom{r-1}{s-1}$.
Proof of the claim.
Fix a vertex $x$ in $R$. There are at most $r^{s-2}m_{2}^{s}\leq(\mu/4)m_{2}^{s}\binom{r-1}{s-1}$ edges of $G$, which have one vertex in the part of $x$ and at least two vertices in some part. Put differently, there are at least $(1-1/s^{2}+3\mu/4)m_{2}^{s}\binom{r-1}{s-1}$ edges in $G$, which are $\mathcal{U}$-partite and have a vertex in $x$. Every edge of $R$ incident to $x$ can host at most $m_{2}^{s}$ of these edges. Every non-edge of $R$ incident to $x$ can host at most $4\nu m_{2}^{s}$ of these edges. Thus $m_{2}^{s}\deg_{R}(x)+4\nu m_{2}^{s}\binom{r-1}{s-1}\geq(1-1/s^{2}+3\mu/4)m_{2}^{s}\binom{r-1}{s-1}.$ Solving for $\deg_{R}(x)$ and recalling the choice of constant yields the desired bound. ∎
Since $r$ is divisible by $s$, the claim and Lemma B.2 imply that there is a perfect matching $\mathcal{M}$ in $R$. We will construct a $\mathcal{Q}_{2}$-tiling by finding tuples inside each edge $X\in\mathcal{M}$, as follows. Each edge $X$ of $\mathcal{M}$ corresponds to an $s$-partite family $\{U_{1},\dotsc,U_{s}\}$. We construct a $\mathcal{Q}_{2}$-tiling in $X$ greedily, by finding one tuple after another. Note that as long as the number of used vertices in each part $U_{i}$ is at most $3\nu m_{2}$, there remain at least $\nu m_{2}^{s}\geq dm_{2}^{s}$ edges among the leftover vertices. (Here we used that $d\leq\mu/16=\nu$.) Within the unused vertices in $X$, select subsets $U^{\prime}_{i}\subseteq U_{i}$, for all $1\leq i\leq s$, of size $\beta m_{2}$ each such that $d_{P}(U^{\prime}_{1},\dotsc,U^{\prime}_{s})\geq d$. (This can be done using an averaging argument). We apply Lemma B.1 with $\beta m_{2}$ playing the rôle of $m$, to obtain an $(\varepsilon,d)$-lower-regular $\tilde{m}$-balanced $s$-tuple, where $m_{3}=2\gamma\beta m_{2}\leq\tilde{m}\leq\beta m_{2}$. Note that this lower-regular $s$-tuple belongs to $\mathcal{Q}_{2}$. We iterate this, to greedily find a $\mathcal{Q}_{2}$-tiling on at least $3\nu sm_{2}$ vertices, and at most $3\nu sm_{2}+\beta m_{2}$ vertices, in the $X$-partite subgraph of $P$ hosted by $\bigcup X$.
Denote the union of these ‘fresh’ $\mathcal{Q}_{2}$-tilings, over all $X\in\mathcal{M}$, by $\mathcal{Q}_{\text{fresh}}$. Note that $\mathcal{Q}_{\text{fresh}}$ covers at least $3\nu m_{2}$ and at most $(3\nu+\beta)m_{2}$ vertices in each part of $\mathcal{U}$. Since there are at least $(1-\lambda){(n-sm_{2})}/m_{2}$ parts in $\mathcal{U}$ outside $V(Q_{1})$, it follows that $\mathcal{Q}_{\text{fresh}}$ covers at least ${3\nu(1-\lambda)n-s}\geq 2\nu\eta n=4\beta n$ vertices outside $V(Q_{1})$.
Next, we define a $\mathcal{Q}_{2}$-tiling $\mathcal{Q}_{\text{rec}}\subseteq P$ in $V(Q_{1})\smallsetminus V(\mathcal{Q}_{\text{fresh}})$ to ‘recycle’ what is left of $Q_{1}$. Let $\{V_{1},\dotsc,V_{s}\}$ be the parts of an ${\tilde{m}}$-balanced $(\varepsilon,d)$-lower-regular $s$-tuple in $Q_{1}$. Recall that, $\mathcal{Q}_{\text{fresh}}$ covers between $3\nu m_{2}$ and $(3\nu+\beta)m_{2}$ vertices in each part of $\mathcal{U}$. This implies in particular that the difference of leftover vertices between any two clusters of $V_{1},\dots,V_{s}$ is at most $\beta{\tilde{m}}$. We may therefore pick subsets $U_{1}\subseteq V_{1},\dotsc,U_{s}\subseteq V_{s}$ of size $\eta{\tilde{m}}$ of size $m^{\ast}=(1-3\nu-\beta){\tilde{m}}$ among the uncovered vertices. Note that $(U_{1},\dots,U_{s})$ is still $(2\varepsilon,d^{\prime})$-lower-regular for $d^{\prime}\geq d-\varepsilon\geq d/2$ and thus in $\mathcal{Q}_{2}$. Indeed, for all choices of $X_{1}\subseteq U_{1},$ $\dots$,… , $X_{s}\subseteq U_{s}$ of size at least $2\varepsilon m^{\ast}\geq\tilde{m}$ each, we have $d_{P}(X_{1},\dots,X_{s})\geq d-\varepsilon$ by $(\varepsilon,d)$-regularity of $(V_{1},\dots,V_{s})$. Let $\mathcal{Q}_{\text{rec}}\subseteq P$ be the $\mathcal{Q}_{2}$-tiling obtained this way. It follows that $\mathcal{Q}_{\text{rec}}\cup\mathcal{Q}_{\text{fresh}}$ covers all but $\beta v(Q_{1})=\beta\lambda n$ vertices of $V(Q_{1})$. It follows that $\mathcal{Q}_{\text{fresh}}\cup\mathcal{Q}_{\text{rec}}$ covers at least $(1-\beta)\lambda n+4\beta n\geq(\lambda+4\beta-2\beta)n=(\lambda+\nu\eta)n$ vertices. ∎